OpenStudy (jkbo):

Please help What is the solution of the matrix equation? [6 5] [x] [2] [5 4] [y] = [1] A.(-2, 1) B.(10, 6) C.(–4, 3) D.(–3, 4)

OpenStudy (anonymous):

what class is this for?

OpenStudy (jkbo):

Alg. 2

OpenStudy (anonymous):

so does row reduction mean anything to you? I'm guessing not, but that's ok. There's an easier way.

OpenStudy (jkbo):

No, it doesn't I just learned this

OpenStudy (anonymous):

just a second I'm double checking my answer so I don't lie to you :)

OpenStudy (anonymous):

ok so when we multiply a matrix like this, we start at the top left and multiply whatever is there by the thing on top of the next little matrix. Then we move over and multiply the next number to the right by the second one down in the second matrix. Here's an animation: http://www.facstaff.bucknell.edu/mastascu/elessonsHTML/Circuit/MatVecMultiply.htm

OpenStudy (anonymous):

you may have to scroll down just a little to see it.

OpenStudy (anonymous):

so the equations we're going to set up is going to look like this: \[6x+5y=2\]and\[5x+4y=1\]

OpenStudy (anonymous):

the tricky thing is getting the right numbers paired up together. Remember, go across the row in the first guy and down the column in the second guy.

OpenStudy (anonymous):

then you just need to solve those equations for x and y by whatever method you prefer. What answer did you get?

OpenStudy (jkbo):

I got c. but I dont think thats right

OpenStudy (anonymous):

It's not, but it's awfully close!

OpenStudy (anonymous):

If we solve the second eq for x, then \[x=\frac{ 1 }{ 5 }-\frac{ 4 }{ 5}y\]

OpenStudy (anonymous):

oh man I picked the hard way. Subbing that in for x in the first one gives\[\frac{ 6 }{ 5}+\frac{ 24 }{ 5 }y+5y=2\]

OpenStudy (anonymous):

whoops the 24 should be negative

OpenStudy (anonymous):

\[\frac{ 1 }{ 5 }y=\frac{ 4 }{ 5}\]

OpenStudy (anonymous):

so y=4

OpenStudy (jkbo):

o ok than its d

OpenStudy (anonymous):

right. do you understand the pattern?

OpenStudy (jkbo):

yes, a little bit

OpenStudy (anonymous):

did you watch that animation? That's probably the best way I can explain it over the internet.

OpenStudy (jkbo):

yes , I did it was helpful